OpenStudy (anonymous):

Solve x2 + 2x + 9 = 0.

OpenStudy (anonymous):

@pooja195

Directrix (directrix):

Is this x^2 + 2x + 9 = 0?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

hope you are allowed complex solutions yes?

OpenStudy (anonymous):

because the two answers will not be real numbers

OpenStudy (anonymous):

Suppose so. These are my choices. http://prntscr.com/9262wo

OpenStudy (anonymous):

oh ok since 2 is even, easiest to complete the square

OpenStudy (anonymous):

subtract 9 from both sides what do you get?

OpenStudy (anonymous):

4.5

OpenStudy (anonymous):

Sorry I'm a little disoriented.

OpenStudy (anonymous):

\[x^2+2x+9=0\] subract \(9\) from both sides

OpenStudy (anonymous):

x^2 + 2x = -9

OpenStudy (anonymous):

ok good

OpenStudy (anonymous):

now what is half of 2?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

right and \(1^2=1\) so go right to \[(x+1)^2=-9+1\\ (x+1)^2=-8\]

OpenStudy (anonymous):

that means \[x+1=\pm\sqrt8\\ z=-1\pm\sqrt{-8}\]

OpenStudy (anonymous):

Is there more to do?

OpenStudy (anonymous):

write \(\sqrt{-8}\) as \(\sqrt{8}i\)

OpenStudy (anonymous):

subtract 1 from both sides?

OpenStudy (anonymous):

then since \(\sqrt8=2\sqrt2\) final answer is \[-1\pm2\sqrt2 i\]

OpenStudy (anonymous):

Oh!

OpenStudy (anonymous):

Wouldn't it be -1 +2i sqrt 2?

OpenStudy (anonymous):

http://prntscr.com/9264c1 The answer choices put the i after the 2.

OpenStudy (anonymous):

yeah what you said

OpenStudy (anonymous):

Alright, thanks.